OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line that passes through (-3,1) and is parallel to to the line represented by y=-2/3x - 5.

OpenStudy (anonymous):

yes

OpenStudy (candyme):

So slope-intercept form is y=mx+b. if your line needs to cross through (-3,1) and be parallel to y=-2/3x-5 you need to set up an equation. Start with the parallel part. If two lines are parallel, they MUST have the same slope. So what would your slope be?

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

what about this? Write the slope-intercept form of the equation of the line that passes through (4,2) an is parallel to the line represented by -2x+ y=7

OpenStudy (candyme):

Did you finish the first one?

OpenStudy (anonymous):

yes

OpenStudy (candyme):

You do the same thing for this one, but you have to get the original equation into slope intercept form. It would be y=2x+7. Since they have to have the same slope to be parallel, your new equation would be y=2x+b Now plug in your point. 2=2(4)+b multiply 2=8+b subtract the 8 -6=b so your equation is y=2x-6

OpenStudy (anonymous):

thank you so much!

OpenStudy (candyme):

No problem glad I could help